Samuel leaves the end of a 4.0-m-high diving board and strikes the 

water 1.3 s later, 3.0 m beyond the end of the board. Considering the 

diver as a particle, determine: 

(a) His initial velocity,

 (b) The maximum height reached;

 (c) The velocity Vf

 with which he enters the water.


1
Expert's answer
2021-08-31T08:58:15-0400

Given:

"H=4.0\:\rm m"

"L=3.0\:\rm m"

"\tau=1.3\:\rm s"


The equations of motion of a diver

"x(t)=v_0\cos\theta *t\\
y(t)=H+v_0\sin\theta*t-\frac{gt^2}{2}"

We get:

"L=v_0\cos\theta *\tau\\
0=H+v_0\sin\theta*\tau-\frac{g\tau^2}{2}"

"3=v_0\cos\theta *1.3\\
0=4+v_0\sin\theta*1.3-\frac{9.8*1.3^2}{2}"

"v_0\cos\theta=2.3\\
v_0\sin\theta=3.3"

(a)

"v_0=\sqrt{2.3^2+3.3^2}=4.0\:\rm m/s""\sin\theta=3.3/4.0=0.82"

(b)

"v_0\sin\theta-gt^*=0\\
t^*=3.3/9.8=0.34\:\rm s"

"H_{\rm max}=y(t^*)=H+v_0\sin\theta*t^*-\frac{gt^{*2}}{2}\\
=4+3.3*0.34-\frac{9.8*0.34^2}{2}=4.6\:\rm m"

(c)

"v_f=\sqrt{v_0^2+2gH}=\sqrt{4.0^2+2*9.8*4}=9.7\:\rm m/s"